# supplemental

### v2

@@ 3.4 Training and Inference
We finetune the pre-trained DiT model using the AdamW optimizer on 8 Nvidia A100 GPUs. The learning rate is set at $0.99\times 10^{-4}$ with an effective batch size of 64. The model is finetuned on $256\times 256$ resolution for about 70K steps, which takes 90 hours. During training, we feed data from our two training datasets Scenes and Materials in a 5:3 ratio, prioritizing the task of image-conditioned material generation. We also randomly drop the conditions to retain the capacity to use classifier-free guidance (CFG) [22]. For text-only or unconditional generation, the mask is replaced by a completely white image placeholder.
@@
Our model completes a generation in 12 seconds using 50 diffusion steps. The model natively outputs a resolution of 256 due to limited computational resources. We apply an upsampler [32] to increase the resolution of each material map to 512 × 512.
4 Results
@@
We evaluate the performance of our MaterialPicker across multiple dimensions. First, we perform qualitative and quantitative comparisons with Material Palette [28] on material extraction using both synthetic and real-world images (Sec. 4.2). Next, we compare with a texture rectification method on real-world images (Sec. 4.2) and with MatGen [47] and MatFuse [48] on text-to-material generation (Sec. 4.3). Finally, we conduct ablation studies on multi-modality, dataset design and evaluate the impact of the input image scale(Sec. 4.4). We also ablate the usage of a mask and evaluate robustness to distortion and lighting/shadowing in the supplemental materials.
4.1 Evaluation dataset and metrics
@@
Synthetic Evaluation dataset. For systematic evaluation, we build a synthetic evaluation dataset by gathering a diverse set of 531 materials from PolyHaven [36], applied to three interior scenes from the Archinteriors collection [14] (which are completely independent from our training set). For each scene, we sequentially apply the 531 collected materials to a designated object inside the scene, and render 2D images using Blender Cycles [4] with the scene’s default illumination setup. We generate a total of 1,593 synthetic renderings, and crop a square around the location of the object with replaced materials.
@@
Real Photographs Evaluation Dataset. To validate the generalization of our models, we curate an evaluation dataset containing real photographs captured by smartphones. This dataset covers a comprehensive set of real-world materials observed under both natural outdoor lighting and complex indoor illumination. We crop the photographs with a primary focus on our target material, without strictly limiting the cropping boundaries.
@@
Evaluation Metrics. Since we do not target pixel-aligned material capture, per-pixel metrics cannot be used for our results. Instead, we focus on the appearance similarity of the materials extracted from the photo inputs. Following related work on high-fidelity image synthesis such as DreamBooth [40], we leverage CLIP-I, which is the average pairwise cosine similarity between ViT-L-14 CLIP [37] embeddings of two sets of images. We also use the DINO metric [40] to measure the average pairwise cosine similarity between ViT-L-16 DINO embeddings. Additionally, we report the FID score [21] to measure the statistical visual similarity between two image distributions. We compute the FID score for each of the 531 output material map sets against the corresponding ground truth material map sets, and average the FID scores over the three scenes in our synthetic evaluation dataset.
4.2 Image Conditioned Generation.
@@
Figure 3: Comparisons with Material palette [28] on synthetic dataset for material extraction. The first column shows the ground truth material maps from PolyHaven, with the rendered scene below. The yellow square area indicates the crop used as the input for both models. The second and third columns show the material maps extracted by our model and Material Palette, along with the re-rendered images. We can see that our approach better matches the Ground-Truth appearance.
We evaluate the performance of our model on both synthetic images and real photographs. We first show a visual comparison with the state-of-the-art method Material Palette [28] on our synthetic evaluation dataset (Sec. 4.1). Since Material Palette generates only three material maps (albedo, normal, and roughness), we present both qualitative and quantitative results for these channels, along with the re-rendered images using these generated material maps. Our method takes 12 seconds to generate a material while Material Palette takes 3 minutes, on the same Nvidia A100 GPU, a 15 times speedup. Furthermore, our model can generate materials in batches. In Fig. 3 we show that our model produces material maps with a closer texture appearance and better matching the ground-truth material maps. In contrast, Material Palette struggles to reconstruct structured textures often resulting in distorted lines. We also observe that in the rendered images, our generated materials better matches the original input images.
We include a quantitative comparison with Material Palette on the entire synthetic dataset in Tab. 1. We find that our proposed model performs better on all three metrics for the vast majority of the generated material channels and the corresponding re-rendered images.
Table 1: Quantitative results of material extraction. We compare with Material Palette [28] and report the average CLIP-I metric $\uparrow$ and DINO metric $\uparrow$ between the output material maps and ground truth. We also report the FID metric $\downarrow$ between the generated and ground-truth material maps. The 95% confidence interval can be found in the supplemental materials.
@@
Mat-Palette 0.503 0.631 0.502 0.797
Ours 0.494 0.672 0.566 0.863
[…]
Mat-Palette 112.249 92.620 155.354 26.828
Ours 107.310 91.057 79.785 12.224
@@
Figure 4: Comparison between our method and Material Palette [28] on material extraction for real photographs. First column shows the input images and the generated (ours bottom-right)/provided (Material Palette top-left) mask for each method. Second to the sixth columns show the generated material maps and rendering under an environment map. We see that our approach better corrects for distortion and match the original appearance.
We show qualitative evaluation on real photographs in Fig. 4 where we can see that our model generalizes well to photographs of materials from various angles. We render the generated materials on a planar surface under environment lighting, showing strong visual similarity to the original input images. Unlike Material Palette, which requires input masks from a separate segmentation step [41], our model operates out-of-box with an input image only, showcasing its potential as a lightweight MaterialPicker.
Since our model automatically performs perspective rectification on the generated materials, we further compare against another state-of-the-art texture rectification and synthesis method [19]. In Fig. 5, we evaluate both methods using real photographs. Since our model directly outputs material maps, instead of textures, we present our results by rendering them under different environment maps. We find that the compared method doesn’t generalize well to real-world photographs, taken from non-frontal and/or non-parallel setups and fails to correct distortion in these cases. In contrast, our approach synthesizes a fronto-parallel view and remains robust across various real-world lighting conditions and viewing angles. Finally, as previously, our model does not require detailed masks as input, directly rectifying the dominant texture in the input image.
@@
(b) Mask (Input)
(c) [Hao et al.]
(d) Our Mask (Output)
(e) Ours (Render 1)
(f) Ours (Render 2)
Figure 5: Comparison with Hao et al. [19] on texture rectification for real photos. The first column shows the input photos. The second and third columns are the required input masks and output textures of Hao et al. The fourth column shows masks generated by our model, followed by two renderings (Render 1 & Render 2) of our generated material maps under different environment maps. Despite not requiring an input mask, our method better rectifies perspective and distortions. Further, as we support material extraction, our result does not contain shading from the input image.
4.3 Text Conditioned Generation.
Although the primary focus of our method is the generation of materials from photos, our multi-modal model also supports text-conditioned generation without image inputs. We evaluate its performance on the text-to-material task, comparing it with two state-of-the-art diffusion-based generative models for material synthesis: MatFuse [48] and MatGen [47]. As shown in Fig. 6, our model demonstrates strong text-to-material synthesis capability, producing high-quality material samples, comparable to other state-of-the-art approaches. Leveraging a pretrained text-to-video model as a prior, our model can interpret complex semantics beyond the material-only training set, such as ”wood rings” and ”floral” patterns.
@@
Figure 6: Comparison of text-to-material generation between our model, MatGen [47], and MatFuse [48]. The ”Text” column contains the input text conditions. The second to last column show the generated material maps, along with a rendering under environment lighting. We note that MatFuse generates a specular map rather than a metallic map.
4.4 Ablation Study
4.4.1 Multi-modality
@@
Figure 7: Comparison of different input combinations. The first column shows the input condition. Text conditioning provides only coarse guidance for material generation, while image inputs offer explicit cues on material appearance. However, image inputs remain ambiguous with respect to material properties, as seen in the third example. Using both text and image conditioning simultaneously reduces this ambiguity, enhancing control and quality.
4.4.2 Mixed Dataset
@@
Figure 8: Impact of our text-to-material synthetic dataset on generation and generalization. For each sample, the first row shows the generation results from our baseline model, trained on mixed datasets (Sec. 3.2), and the second row shows results from a model trained only on the Synthetic scenes. The model trained with mixed dataset is able to synthesize better non-stationary, realistic textures. The leftmost side of each row is labeled with the text conditioning input used.
4.4.3 Input scale
Reproducing the texture scale in the input photos is critical for material generation. As we process our training data to align the scales of input images and output material maps (Sec. 3.2), our model generates scale-matched materials, as shown in Fig. 9. We see that our result follow the scale of the input as it increases from top to bottom.
”Marble tiles”
(a) Input
(b) Albedo
(c) Normal
(d) Roughness
(e) Height
(f) Metallic
(g) Render
”Marble tiles”
”Marble tiles”
Figure 9: Evaluation of our model’s adaptability to different input texture scales on real photographs. We can see that our results are generated with a scale matching that of the input.
5 Limitations
Despite strong generation capacity, our model may still encounter challenging inputs, as shown in Fig. 10. In the first row we show an example where our model confuses shading and albedo variation. Our model may also have difficulty handling materials with cutouts or holes, since it does not produce opacity maps as outputs. Finally, preserving semantically meaningful patterns, such as text, is a remaining challenge in our approach.
“Concrete
pavement”
(a) Input
(b) Albedo
(c) Normal
(d) Roughness
(e) Height
(f) Metallic
(g) Render
“Perforated
metal”
“Towel”
Figure 10: Limitations. We show limitations of our model, such as complex lighting and shadows in the first row, materials with perforations in the second row, and structurally significant elements like text in the third row.
6 Conclusion
We present a generative model for high-quality material synthesis from text prompts and/or crops of natural images by finetuning a pretrained text-to-video generative model, which provides strong prior knowledge. The flexible video DiT architecture lets us adjust the model for multi-channel material generation. We show extensive evaluation on both synthetic and real examples and conduct systematic ablation studies and test on robustness. We believe that our re-purposing of a video model for multi-channel generation opens an interesting avenue for other domain which require the generation of additional channels, such as intrinsic decomposition [47].
Supplementary Material
Appendix A More Results
A.1 Image Conditioned Generation
We show qualitative and quantitative comparisons with Material Palette [28] on material extraction in Sec. 4.2 of the main paper. We provide more visual comparisons in Fig. 15 and Fig. 16. Since Material Palette generates only three material maps (albedo, normal, and roughness), we present results for these channels, along with the re-rendered images. We report the average CLIP-I metric $\uparrow$ and DINO metric $\uparrow$ between the output material maps and ground truth in Tab. 1 in the main paper. We further report the 95% confidence interval in Tab. 2 of the Supplemental Material. Our method achieves higher 95% confidence intervals for the vast majority of the generated material channels with the exception of the Albedo for which the intervals overlap. Our re-rendered images also show consistently higher alignment with the ground truth.
Table 2: Quantitative results of material extraction. We compare with Material Palette [28] and present the 95% confidence intervals for the CLIP-I metric $\uparrow$ and the DINO metric $\uparrow$, calculated between the output material maps and the ground truths.
[…]
Mat-Palette (0.814, 0.820) (0.864, 0.870) (0.788, 0.794) (0.954, 0.956)
Ours (0.855, 0.860) (0.872, 0.876) (0.863, 0.869) (0.966, 0.968)
[…]
Mat-Palette (0.494, 0.513) (0.622, 0.640) (0.493, 0.512) (0.792, 0.802)
Ours (0.484, 0.503) (0.663, 0.680) (0.556, 0.576) (0.859, 0.867)
A.2 Ablation Study
A.2.1 Mask as Input or Output
@@
As opposed to existing material generation models, our model doesn’t require the target material to cover the entire input image [47] or manually-created masks [28] to identify the sample of interest. Our model instead outputs a mask along with the generated materials. To assess the impact of generating this mask, we train an alternative model using our two datasets, with a slight modification to the model configuration. We add noise $\epsilon_{t}$ to the material maps $M$ only, with $x_{t}=\text{stack}(I,V,M_{t})$, leaving the image and mask as non-noised inputs (or $x_{t}=\text{stack}(V,M_{t})$ without $I$), using our adaptation of a video model (as described in Sec. 3.3). The loss is then computed on the material maps $M$ only. As shown in Fig. 11, we find that our proposed model, which automatically predicts a mask, performs comparably well to this variant requiring the mask as input.
”Ceramic
tiles”
@@
Figure 11: Impact of mask on material generation quality. Here shows the effect of using the mask as an input versus as an output on the quality of generated materials. Each pair of consecutive rows represents the results from the model with the mask as input (top row) and the model with the mask as output (bottom row). The results show that our model can accurately predict masks without a decrease in material quality. The leftmost side of each row is labeled with the text conditioning input used.
A.3 Evaluations on the Robustness.
We further test our model’s robustness under various factors, including different texture scales (in the main paper), varying levels of distortion, and diverse lighting conditions.
To examine the robustness of our model to strong, real-world, distortions, we generate a synthetic test set that use textures from the texture datasets TexSD [28] and follow the texture processing steps outlined by Hao et al. [19]. We apply homography transformations [20] and thin plate spline transformations [5] to the textures. Our results in Fig. 12 show that the model is robust to severe distortions, stretching, and the blurring effects introduced by these transformations. More examples of real photos with distortion or surface geometry diversity can be found in Fig. 17, Fig. 18, Fig. 19, Fig. 20, Fig. 21 and Fig. 22.
”Gingham”

### v3

@@ 3.4. Training and Inference
We finetune the pre-trained DiT model using the AdamW optimizer on 8 Nvidia A100 GPUs. The learning rate is set at $0.99\times 10^{-4}$ with an effective batch size of 64. The model is finetuned on $\times 256$ 256 resolution for about 70K steps, which takes 90 hours. During training, we feed data from our two training datasets Scenes and Materials in a 5:3 ratio, prioritizing the task of image-conditioned material generation. For text-only or unconditional generation, the mask is replaced by a completely white image placeholder.
@@
Our model completes a generation in 12 seconds using DDIM (Song et al., 2020) with 50 diffusion steps on an Nvidia A100. The model natively outputs a resolution of 256 due to limited computational resources. We apply an upsampler (Niu, 2023) to increase the resolution of each material map to 512 $\times$ 512.
4. Results
@@
We evaluate the performance of our MaterialPicker across multiple dimensions. First, we perform qualitative and quantitative comparisons with Material Palette (Lopes et al., 2024) on material extraction using both synthetic and real-world images (Sec. 4.2). Next, we compare with a material acquisition method (Vecchio et al., 2024a) and a texture rectification method on real-world images (Sec. 4.2), and with MatGen (Vecchio et al., 2024a) and MatFuse (Vecchio et al., 2024b) on text-to-material generation (Sec. 4.3). Finally, we conduct ablation studies on multi-modality, dataset design, the usage of a mask, and evaluate the impact of the input image scale and the robustness to distortion and lighting/shadowing. We also demonstrate our generalizability to complex patterns and the ability to generate seamless results (Sec. 4.4).
4.1. Evaluation dataset and metrics
4.1.1. Synthetic evaluation dataset.
@@ 4.1. Evaluation dataset and metrics
For systematic evaluation, we build a synthetic evaluation dataset by gathering a diverse set of 531 materials from PolyHaven [^1] https://polyhaven.com/, applied to three interior scenes from the Archinteriors collection (Evermotion, 2021) (completely independent from our training set). For each scene, we sequentially apply the 531 collected materials to a designated object inside the scene, and render 2D images using Blender Cycles (Blender Community, 2018) with the scene’s default illumination setup. We generate 1,593 synthetic renderings, and crop a square around the location of the object with replaced materials.
4.1.2. Real photographs evaluation dataset.
@@
To validate the generalization of our models, we curate an evaluation dataset containing real photographs captured by smartphones. This dataset covers a comprehensive set of real-world materials observed under both natural outdoor lighting and complex indoor illumination. We crop the photographs with a primary focus on our target material, without strictly limiting the cropping boundaries.
4.1.3. Evaluation metrics.
@@
Since we do not target pixel-aligned material capture, per-pixel metrics cannot be used for our results. Instead, we focus on the appearance similarity of the materials extracted from the photo inputs. Following related work on high-fidelity image synthesis such as DreamBooth (Ruiz et al., 2023), we leverage CLIP-I, which is the average pairwise cosine similarity between ViT-L-14 CLIP (Radford et al., 2021) embeddings of two sets of images. We also use the DINO metric (Ruiz et al., 2023) to measure the average pairwise cosine similarity between ViT-L-16 DINO embeddings.
4.2. Image Conditioned Generation
@@
Figure 3. Comparisons with Material Palette (Lopes et al., 2024) on synthetic dataset for material extraction. The first column shows the ground truth material maps from PolyHaven, with the rendered scene below. The yellow square area indicates the crop used as the input for both models. The second and third columns show the material maps extracted by our model and Material Palette, along with the re-rendered images. We can see that our approach better matches the ground truth appearance.
We evaluate the performance of our model on both synthetic images and real photographs. We first show a visual comparison with the state-of-the-art method Material Palette (Lopes et al., 2024) on our synthetic evaluation dataset (Sec. 4.1.1). Since Material Palette generates only three material maps (albedo, normal, and roughness), we present both qualitative and quantitative results for these channels, along with the re-rendered images using these generated material maps. Our method takes 12 seconds to generate a material while Material Palette takes 3 minutes, on the same Nvidia A100 GPU, a 15 times speedup. Furthermore, our model can generate materials in batches. In Fig. 3 we show that our model produces material maps with a closer texture appearance and better matching the ground-truth material maps. In contrast, Material Palette struggles to reconstruct structured textures often resulting in distorted lines. We also observe that in the rendered images, our generated materials better matches the original input images.
We include a quantitative comparison and 95% confidence interval with Material Palette on the entire synthetic dataset in Tab. 1. We find that our model performs better on two metrics for the vast majority of generated materials channels, with the exception of the Albedo for which the intervals overlap. Our re-rendered images also show consistently higher alignment with the ground truth.
Table 1. Quantitative results of material extraction. We compare with Material Palette (Lopes et al., 2024) and report the average CLIP-I metric $\uparrow$ and DINO metric $\uparrow$ between the output material maps and ground truth alongside the 95% confidence interval.
@@
Mat-Palette 0.503±0.1$\pm$ 0.1± 0.1 0.631±0.09$\pm$ 0.09± 0.09 0.502±0.09$\pm$ 0.09± 0.09 0.797±0.05$\pm$ 0.05± 0.05
Ours 0.494±0.1$\pm$ 0.1± 0.1 0.672±0.08$\pm$ 0.08± 0.08 0.566±0.1$\pm$ 0.1± 0.1 0.863±0.04$\pm$ 0.04± 0.04
@@
Figure 4. Comparison between our method and Material Palette (Lopes et al., 2024) on material extraction for real photographs. The first column shows the input images and the generated (ours bottom-right)/provided (Material Palette top-left) masks. The second to sixth columns show the generated material maps and rendering under an environment map. We see that our approach better corrects for distortion and match the original appearance.
We show qualitative evaluation on real photographs in Fig. 4 where we see that our model generalizes well to photographs of materials from various angles. We render the generated materials on a planar surface under environment lighting, showing strong visual similarity to the original input images. Unlike Material Palette, which requires input masks from a separate segmentation step (Sharma et al., 2023), our model operates out-of-box with an input image only, showcasing its potential as a lightweight MaterialPicker.
We show a comparison with ControlMat (Vecchio et al., 2024a) in Fig. 5. Since ControlMat uses local features of the images extracted by ControlNet (Zhang et al., 2023) to guide the diffusion process, it is trained to generate results that align with the input, therefore fail to handle imperfect perspectives and distortions. Further, since the generation of multichannel materials in ControlMat depends on modifications to the VAE, it must be trained from scratch without leveraging any image or video priors from which our method benefits. This limits ControlMat generalization to complex textures captured under unseen lighting conditions (e.g. 4th row).
Ours
@@
(b) Albedo
(c) Normal
(d) Roughness
(e) Height
(f) Metallic
(g) Render
ControlMat
Ours
ControlMat
Ours
ControlMat
Ours
ControlMat
Figure 5. Comparison between our method and ControlMat (Vecchio et al., 2024a) on material extraction for real photographs with distortions. The first column shows the input images. The second to the sixth columns show the generated material maps and rendering under an environment map. As ControlMat relies on well aligned conditions through ControlNet, our approach demonstrates superior performance in correcting perspectives and distortions. Further, as ControlMat is trained only on material data, it does not benefit from the learned priors from a pre-trained model like our fine-tuning, making their method less effective at generalizing to unseen patterns and lighting conditions (e.g. in the 4th row).
Since our model automatically performs perspective rectification on the generated materials, we further compare against another state-of-the-art texture rectification and synthesis method (Hao et al., 2023). In Fig. 17, we evaluate both methods using real photographs. Since our model directly outputs material maps, instead of textures, we present our results by rendering them under different environment maps. We find that the compared method does not generalize well to real-world photographs, taken from non-frontal and/or non-parallel setups and fails to correct distortion in these cases. In contrast, our approach synthesizes a fronto-parallel view and remains robust across various real-world lighting conditions and viewing angles. Finally, as previously, our model does not require detailed masks as input, directly rectifying the dominant texture in the input image.
4.3. Text Conditioned Generation
Although the primary focus of our method is the generation of materials from photos, our multi-modal model also supports text-conditioned generation without image inputs. We evaluate its performance on the text-to-material task, comparing it with two state-of-the-art diffusion-based generative models for material synthesis: MatFuse (Vecchio et al., 2024b) and MatGen (Vecchio et al., 2024a). As shown in Fig. 6, our model demonstrates strong text-to-material synthesis capability, producing high-quality material samples, comparable to other state-of-the-art approaches. We report the cosine similarity between each rendered image and the text condition, calculated using ViT-L-14 CLIP (Radford et al., 2021) embeddings. Leveraging a pretrained text-to-video model as a prior, our model can interpret complex semantics beyond the material-only training set, such as "wood rings" and "floral" patterns.
@@ 4.3. Text Conditioned Generation
0.132
Figure 6. Comparison of text-to-material generation between our model, MatGen (Vecchio et al., 2024a), and MatFuse (Vecchio et al., 2024b). The "Text" column contains the input text conditions. The second to last columns show the generated material maps, along with a rendering under environment lighting. The CLIP score between the rendering and the text is reported at the bottom of each rendered image (higher is better). We note that MatFuse generates a specular map rather than a metallic map.
4.4. Ablation Study
4.4.1. Multi-modality
@@
Figure 7. Comparison of different input combinations. The first column shows the input condition. Text conditioning provides only coarse guidance for material generation, while image inputs offer explicit cues on material appearance. However, image inputs remain ambiguous with respect to material properties, as seen in the third example. Using both text and image conditioning simultaneously reduces this ambiguity, enhancing control and quality.
4.4.2. Mixed Dataset
@@
Figure 8. Impact of our text-to-material synthetic dataset on generation and generalization. For each sample, the first row shows the generation results from our baseline model, trained on mixed datasets (Sec. 3.2), and the second row shows results from a model trained only on the Synthetic scenes. The model trained with mixed dataset is able to synthesize better non-stationary, realistic textures. The leftmost side of each row is labeled with the text conditioning input used.
4.4.3. Mask as Input or Output
@@
As opposed to existing material generation models, our model does not require the target material to cover the entire input image (Vecchio et al., 2024a) or manually-created masks (Lopes et al., 2024) to identify the sample of interest. Our model instead outputs a mask along with the generated materials. To assess the impact of generating this mask, we train an alternative model using our two datasets, with a slight modification to the model configuration. We add noise $\epsilon_{t}$ to the material maps $M$ only, with $x_{t}=\text{stack}(I,V,M_{t})$, leaving the image and mask as non-noised inputs (or $x_{t}=\text{stack}(V,M_{t})$ without $I$), using our adaptation of a video model (as described in Sec. 3.3). The loss is then computed on the material maps $M$ only. As shown in Fig. 9, we find that our proposed model, which automatically predicts a mask, performs comparably well to this variant requiring the mask as input.
"Ceramic
tiles"
@@
Figure 9. Impact of mask on material generation quality. Here shows the effect of using the mask as an input versus as an output on the quality of generated materials. Each pair of consecutive rows represents the results from the model with the mask as input (top row) and the model with the mask as output (bottom row). The results show that our model can accurately predict masks without a decrease in material quality. The leftmost side of each row is labeled with the text conditioning input used.
4.4.4. Input scale
Reproducing the texture scale in the input photos is critical for material generation. As we process our training data to align the scales of input images and output material maps (Sec. 3.2), our model generates scale-matched materials, as shown in Fig. 10. We see that our result follows the scale of the input as it increases from top to bottom.
"Marble tiles"